# Impact of correlations on nuclear binding energies

### v1

@@ Abstract
A strong effort will be dedicated in the coming years to extend the reach of ab initio nuclear-structure calculations to heavy doubly open-shell nuclei. In order to do so, the most efficient strategies to incorporate dominant many-body correlations at play in such nuclei must be identified. With this motivation in mind, the present work pedagogically analyses the inclusion of many-body correlations and their impact on binding energies of Calcium and Chromium isotopes.
@@
Eventually, the present work demonstrates in a pedagogical way that polynomially-scaling expansion methods based on unperturbed states that possibly break (and restore) symmetries constitute an optimal route to extend ab initio calculations to heavy closed- and open-shell nuclei.
@@ 1 Introduction
In this context, the present work wishes to pedagogically analyse the impact of many-body correlations on binding energies and associated differential quantities, i.e. first- and second-order derivatives with respect to the (even) neutron number. To do so, the study is conducted along neighboring Calcium ($Z=20$) and Chromium ($Z=24$) isotopic chains spanning a large range of (even) neutron numbers from $N=12$ till $N=50$. Most of Ca isotopes are of singly open-shell character whereas most of Cr isotopes are of doubly open-shell character. Comparing the behavior of binding energies along these isotopic chains allows one to illustrate the roles played by static and dynamical correlations in the two types of nuclei and the capacity of ab initio many-body methods to efficiently capture them by employing an optimal formulation. In order to control how some of the identified features depend on the nuclear mass, additional calculations are performed along the Tin ($Z=50$) isotopic chain from $N=50$ till $N=82$.
@@
Figure 1: Systematics along Ca (left) and Cr (right) isotopic chains. First line: sHFB absolute binding energies against experimental data. Second (third) line: sHFB and sHF-EFA two-neutron separation energy (two-neutron shell gap) against experimental data. Between 42Ca and 48Ca (1f7/2 shell) as well as between 62Ca and 70Ca (1g9/2 shell), sHFB-ZP semi-analytical results are also shown. Fourth line: sHFB neutron-number variance against the minimal possible variance in sHFB calculations Duguet et al. (2020). The sequence of the underlying neutron canonical shells are also displayed.
@@ 3.1 Ca chain
The main characteristics of the experimental $S_{2n}$, i.e. the sudden drops at $N=20$ and $28$, and to a lesser extent at $N=32$ and $34$, as well as the smooth evolution in between, are well accounted for by sHFB results. However, crucial differences are revealed upon closer inspection. First, the amplitude of the drops at $N=20$ and $28$ is too large and the trend in between, i.e. while filling the 1f7/2 shell, is qualitatively wrong. Correlated with the too large drop at $N=20$, the $S_{2n}$ value in 42Ca is significantly too low. Further adding neutrons, $S_{2n}$ increases linearly throughout the 1f7/2 shell instead of decreasing linearly as for experimental data [^2] While the same patterns are at play when going through 48Ca, the size of the 2p3/2 shell is too small to make the rising slope of sHFB results really visible. The highly degenerate 1g9/2 shell between 62Ca and 70Ca is more favorable in this respect even though the slope of the sHFB results is actually zero in this case. These nuclei are anyway predicted to be unbound and there is no experimental data yet to be confronted with..
Given that $S_{2n}(N,Z)$ is the first derivative of the binding energy, the patterns identified above relate to specific features of the binding energies that could not be fully appreciated in the first line of Fig. 1 due to the large scale employed. The fact that $S_{2n}$ evolves linearly with the number $a_{v}$ of nucleons in the valence shell for both sHFB results and experiment data implies that $E(N,Z)$ is essentially quadratic in between two closed-shell isotopes. The fact that $S_{2n}$ starts from a too low value in sHFB calculations in the open-shell relates to the fact that the linear decrease of $E(N,Z)$ is not pronounced enough such that the difference to the data increases throughout the shell. Finally, the fact $S_{2n}$ is rising linearly instead of decreasing linearly indicates that the sHFB energy is concave instead of being convex.
Figure 2: Energy curvature of Ca (left) and Cr (right) isotopes between $N=20$ and $N=28$ (1f7/2 shell). Upper panels: energies rescaled to $N=20$ and $N=28$ (see text for details). Bottom panels: two-neutron shell gap $\Delta_{2n}$. Experimental data are compared to sHFB results. For Ca isotopes, results from sHF-EFA and sHFB-ZP semi-analytical results are also shown.
@@ 3.2 Analytical investigation
In order to proceed, one must first make a crucial observation thanks to the results shown on the last line of Fig. 1 comparing the neutron-number variance in the sHFB calculation to the minimal variance obtained in the zero-pairing limit of sHFB theory (sHFB-ZP) Duguet et al. (2020). As already noticed Somà et al. (2021), chiral Hamiltonians typically generate only little static pairing at the mean-field level [^5] In an ab initio setting, pairing properties such as the odd-even mass staggering are expected to largely originate from (i.e. required to account for) higher-order processes associated with the exchange of collective medium fluctuations between paired particles Barranco et al. (2004); Gori et al. (2005); Pastore et al. (2008); Idini et al. (2011). Achieving a quantitative description of pairing properties from first principles constitutes a major challenge for ab initio nuclear structure theory Somà et al. (2021)., i.e. the computed neutron-number variance is indeed very close to the minimal variance in most open-shell isotopes, except in 56,58Ca and for nuclei in the continuum. As visible from the left panels of Fig. 1, this is confirmed by the proximity of sHFB results to those obtained from spherical Hartree-Fock calculations performed within the equal-filling approximation Perez-Martin and Robledo (2008) (sHF-EFA) that do not include pairing correlations by construction. Results are indeed very close overall, except in 56,58Ca (1f5/2 shell) where sHFB better reproduce experimental values for $S_{2n}$ and $\Delta_{2n}$. As for the curvature within open-shells, the left panels of Fig. 2 reveals that the curvature of sHF-EFA results also carries the wrong sign but is such that the concavity is even more pronounced than for sHFB, i.e. the weak pairing correlations present within the 1f7/2 shell in sHFB do improve the situation compared to the case where pairing would indeed be strictly zero.
@@
Equation (3a) proves that the sHF-EFA energy is indeed quadratic [^7] As shown in A, the 3N interaction actually induces the presence of a cubic term in the energy. However, present numerical applications demonstrate that it is negligible for all nuclei under consideration such that it can be dropped altogether in the present discussion. in the number of valence nucleons throughout any given open-shell. The coefficient $\alpha_{\breve{v}}$ of the linear term is nothing but the mean-field single-particle energy of the valence shell computed in the CS core $\epsilon^{\text{CS}}_{\breve{v}}$, whose interaction energy contributions are displayed diagrammatically in Fig. 3. The coefficient $\beta_{\breve{v}}$ of the quadratic term, i.e. the curvature, of the energy is given by the average over the valence magnetic substates of the diagonal valence-shell two-body matrix elements [^8] As seen from Eq. (4b), $\overline{\mathbb{v}}_{vv^{\prime}vv^{\prime}}$ includes the effective contribution obtained by averaging the 3N interaction over the CS core. $\overline{\mathbb{v}}_{vv^{\prime}vv^{\prime}}$ displayed diagrammatically in Fig. 4. Such an averaging corresponds to the monopole valence-shell matrix element per valence state. As visible from Eq. (3b), $-2\epsilon^{\text{CS}}_{\breve{v}}$ sets the initial value of $S_{2n}$ [^9] As seen from Tab. 1, the relation $|\alpha_{\breve{v}}|\gg|\beta_{\breve{v}}|$ holds in practice such that the starting value of $S^{\text{sHF-EFA}}_{2n}$ ($a_{v}=2$) in the open shell is essentially dictated by $\epsilon^{\text{CS}}_{\breve{v}}$. whereas $-2\beta_{\breve{v}}$ drives its linear evolution throughout the open-shell. Eventually, $\Delta_{2n}$ extracts $4\beta_{\breve{v}}$.
@@
Figure 5: Neutron canonical single-particle energies $\epsilon_{\breve{k}}$ from sHFB calculations along the Ca isotopic chain. Semi-analytical sHFB-ZP results for $\epsilon_{1\text{f}_{7/2}}$ are also shown between 40Ca and 48Ca.
Extracting $\epsilon^{\text{CS}}_{\breve{v}}$ and $\overline{\mathbb{v}}_{vv^{\prime}vv^{\prime}}$ numerically from the presently employed chiral Hamiltonian (see Tab. 1), the semi-analytical results from Eqs. (3)-(4) (in fact of their sHFB-ZP counterparts; see A) are superimposed on the left panels of Fig. 1 between 42Ca and 48Ca (1f7/2 valence shell) as well as between 62Ca and 70Ca (1g9/2 valence shell). The results perfectly match the numerical sHF-EFA curves, that are themselves very close to sHFB results. Looking at the left panels of Fig. 2, one indeed sees the fully quantitative agreement between sHF-EFA and the semi-analytical results.
@@
    $E^{\text{sHFB}}$ looses energy relatively to experiment,
@@
    $S^{\text{sHFB}}_{2n}$ starts from a too low value,
@@
the coefficient of the slope being given by $\beta_{\tilde{v}}$. As visible in Fig. 5, neutron canonical single-particle energies do evolve linearly within a given open-shell. In particular, the evolution of $\epsilon_{1\text{f}_{7/2}}$ between 40Ca and 48Ca is perfectly reproduced using Eq. (5) (in fact its sHFB-ZP counterpart; see A). Eventually, this linear down-slopping evolution is fully correlated with the concavity of the binding energy.
@@ 3.3 Cr chain
Having characterized sHFB results along the semi-magic Ca isotopic chain, the focus turns now to doubly open-shell Cr isotopes.
As seen in the upper-right panel of Fig. 1, the global trend of sHFB binding energies is similar, relative to the data, than for Ca isotopes. Looking through the magnifying glass of $S_{2n}$ and $\Delta_{2n}$, experimental data do not however display the characteristic patterns identified along the Ca chain. In particular, $S_{2n}$ decreases more gradually such that the sudden drops (sudden spikes in $\Delta_{2n}$), e.g. at $N=20$ and $28$, have all disappeared. Contrarily, a small bump (spike) is now visible in $S_{2n}$ ($\Delta_{2n}$) for $N=24$, i.e. in 44Cr located in the middle of the 1f7/2 shell. These changes are not at all accounted for by sHFB results that closely follow those obtained previously. Indeed, in addition to displaying the defects identified along the Ca isotopic chain, sHFB results further fail to capture the qualitative modifications seen in the data, i.e. sHFB results keep a strong memory of the underlying spherical shell structure whose fingerprints are no longer visible along the Cr isotopic chain.
@@ 4.1 Ca chain
In spite of the $N=20$ two-neutron shell gap being still overestimated in sBMBPT(2) and sBCCSD, the $S_{2n}$ at the beginning of each open shell is increased to be in much better agreement with experimental data. For example, dynamical correlations bring $S_{2n}$ in 42Ca from $13.3$ MeV in sHFB to $18.0$ and $18.1$ MeV in BMBPT(2) and BCCSD, respectively, as well as to $18.6$ MeV in VS-IMSRG(2), which compares favorably with the experimental value of $19.8$ MeV. Third, the wrong linear increase throughout any given open shell is corrected for, as can be seen for example between 42Ca and 48Ca. This reflects the improvement of the curvature of the energy throughout open shells that can be better appreciated from the left-panels of Fig. 7 that focuses on the 1f7/2 shell. Dynamical correlations turn the energy from beyond concave at the sHFB level to being convex, in a way that is essentially identical with the three employed methods.
Eventually, the agreement with data for $S_{2n}$ and $\Delta_{2n}$ along the Ca chain is qualitatively and quantitatively satisfying for all three methods even though the $N=20$ magicity is still exaggerated in BMBPT(2) and BCCSD and the convexity throughout the 1f7/2 shell is not pronounced enough compared to experimental data for all three methods, which points to yet missing correlations. It will be interesting to investigate in the future whether the lack of convexity in the energy is correlated with the inability of presently employed ab initio methods to correctly reproduce the (infamous) evolution of charge radii between 40Ca and 48Ca Caurier et al. (2001).
@@ 4.2 Analytical investigation
As seen from Eqs. (6)-(8), and in agreement with the results shown in the middle-left panel of Fig. 6 and analyzed in the present section, dynamical correlations modify both the starting value and the slope of $S_{2n}$ in the valence-shell. For example, the negative second-order self-energy correction $\Sigma^{(2)}_{1\text{f}_{7/2}}$ lowers $\epsilon^{\text{CS}(2)}_{1\text{f}_{7/2}}$ in such a way that $S_{2n}$ computed in sBMBPT(2) increases from $13.28$ to $18.02$ MeV in 42Ca to almost match the experimental value ($19.84$ MeV). This effect relates to the coupling of a propagating nucleon to 1-particle/2-hole and 2-particle/1-hole configurations as represented in Fig 8, the latter winning over the former [^11] The lowering of $\epsilon_{1\text{f}_{7/2}}$ is not accompanied by a decrease of $\Delta_{2n}$ in 40Ca in sBMBPT(2) and sBCCSD, contrary to sVS-IMSRG(2), i.e. $\epsilon_{1\text{d}_{3/2}}$ is lowered as much as $\epsilon_{1\text{f}_{7/2}}$. Thus, the needed increase of the effective mass associated with the compression of on-shell single-particle energies is not accounted for by low-order corrections to sHFB.. Consistently, the second-order correction to the average 1f7/2 valence-shell effective interaction is repulsive, with the hole-hole contribution winning over the particle-particle one. In the present calculation, such a correction is larger in absolute value than the mean-field contribution and manages to turn the total energy from being concave to being convex, i.e. it makes $S_{2n}$ decrease linearly between 42Ca and 48Ca as for experimental data [^12] The amount by which $S_{2n}$ is increased at the start of the open-shell and the fact that its slope is actually inverted depend on the Hamiltonian under use; see Refs. Tichai et al. (2018); Somà et al. (2021) for examples where the qualitative defects of the sHFB results are not actually corrected via the inclusion of low-order dynamical correlations.. Still, and as can be seen from the bottom-left panel of Fig. 7, the positive curvature $\beta^{(2)}_{1\text{f}_{7/2}}=25$ keV [^13] This value is essentially constant throughout the valence shell. is not large enough [^14] The same is true for sBCCSD and VS-IMSRG(2) calculations as can be inferred from the bottom-left panel of Fig. 7. compared to experimental data ($\Delta_{2n}/4\approx 220$ keV in 42-46Ca), thus pointing to yet missing many-body correlations as discussed earlier on.

### v2

@@ Abstract
A strong effort will be dedicated in the coming years to extend the reach of ab initio nuclear-structure calculations to heavy doubly open-shell nuclei. In order to do so, the most efficient strategies to incorporate dominant many-body correlations at play in such nuclei must be identified. With this motivation in mind, the present work analyses the step-by-step inclusion of many-body correlations and their impact on binding energies of Calcium and Chromium isotopes.
@@
Eventually, the present work demonstrates that polynomially-scaling expansion methods based on unperturbed states that possibly break (and restore) symmetries constitute an optimal route to extend ab initio calculations to heavy closed- and open-shell nuclei.
@@ 1 Introduction
In this context, the present work wishes to analyse the impact of the step-by-step inclusion of many-body correlations on binding energies and associated differential quantities, i.e. first- and second-order derivatives with respect to the (even) neutron number, while following different possible strategies to do so. The study is conducted along neighboring Calcium ($Z=20$) and Chromium ($Z=24$) isotopic chains spanning a large range of (even) neutron numbers from $N=12$ till $N=50$. Most of Ca isotopes are of singly open-shell character whereas most of Cr isotopes are of doubly open-shell character. Comparing the behavior of binding energies along these isotopic chains allows one to illustrate the roles played by static and dynamical correlations in the two types of nuclei and the capacity of ab initio many-body methods to efficiently capture them by employing an optimal formulation. In order to control how some of the identified features depend on the nuclear mass, additional calculations are performed along the Tin ($Z=50$) isotopic chain from $N=50$ till $N=82$.
@@
Figure 1: Systematics along Ca (left) and Cr (right) isotopic chains. First line: sHFB absolute binding energies against experimental data. Second (third) line: sHFB and sHF-EFA two-neutron separation energy (two-neutron shell gap) against experimental data. Between 42Ca and 48Ca (0f7/2 shell) as well as between 62Ca and 70Ca (0g9/2 shell), sHFB-ZP semi-analytical results are also shown. Fourth line: sHFB neutron-number variance against the minimal possible variance in sHFB calculations Duguet et al. (2020). The sequence of the underlying neutron canonical shells are also displayed.
@@ 3.1 Ca chain
The main characteristics of the experimental $S_{2n}$, i.e. the sudden drops at $N=20$ and $28$, and to a lesser extent at $N=32$ and $34$, as well as the smooth evolution in between, are well accounted for by sHFB results. However, crucial differences are revealed upon closer inspection. First, the amplitude of the drops at $N=20$ and $28$ is too large and the trend in between, i.e. while filling the 0f7/2 shell, is qualitatively wrong. Correlated with the too large drop at $N=20$, the $S_{2n}$ value in 42Ca is significantly too low. Further adding neutrons, $S_{2n}$ increases linearly throughout the 0f7/2 shell instead of decreasing linearly as for experimental data [^2] While the same patterns are at play when going through 48Ca, the size of the 1p3/2 shell is too small to make the rising slope of sHFB results really visible. The highly degenerate 0g9/2 shell between 62Ca and 70Ca is more favorable in this respect even though the slope of the sHFB results is actually zero in this case. These nuclei are anyway predicted to be unbound and there is no experimental data yet to be confronted with..
Given that $S_{2n}(N,Z)$ is the first derivative of the binding energy, the patterns identified above relate to specific features of the binding energies that could not be fully appreciated in the first line of Fig. 1 due to the large scale employed. The fact that $S_{2n}$ evolves linearly with the number $a_{v}$ of nucleons in the valence shell for both sHFB results and experiment data implies that $E(N,Z)$ is essentially quadratic in between two closed-shell isotopes. The fact that $S_{2n}$ starts from a too low value in sHFB calculations in the open-shell relates to the fact that the linear decrease of $E(N,Z)$ is not pronounced enough such that the difference to the data increases throughout the shell. Finally, the fact that $S_{2n}$ is rising linearly instead of decreasing linearly indicates that the sHFB energy is concave instead of being convex.
Figure 2: Energy curvature of Ca (left) and Cr (right) isotopes between $N=20$ and $N=28$ (0f7/2 shell). Upper panels: energies rescaled to $N=20$ and $N=28$ (see text for details). Bottom panels: two-neutron shell gap $\Delta_{2n}$. Experimental data are compared to sHFB results. For Ca isotopes, results from sHF-EFA and sHFB-ZP semi-analytical results are also shown.
@@ 3.2 Analytical investigation
In order to proceed, one must first make a crucial observation thanks to the results shown on the last line of Fig. 1 comparing the neutron-number variance in the sHFB calculation to the minimal variance obtained in the zero-pairing limit of sHFB theory (sHFB-ZP) Duguet et al. (2020). As already noticed Somà et al. (2021), chiral Hamiltonians typically generate only little static pairing at the mean-field level [^5] In an ab initio setting, pairing properties such as the odd-even mass staggering are expected to largely originate from (i.e. required to account for) higher-order processes associated with the exchange of collective medium fluctuations between paired particles Barranco et al. (2004); Gori et al. (2005); Pastore et al. (2008); Idini et al. (2011). Achieving a quantitative description of pairing properties from first principles constitutes a major challenge for ab initio nuclear structure theory Somà et al. (2021)., i.e. the computed neutron-number variance is indeed very close to the minimal variance in most open-shell isotopes, except in 56,58Ca and for nuclei in the continuum. As visible from the left panels of Fig. 1, this is confirmed by the proximity of sHFB results to those obtained from spherical Hartree-Fock calculations performed within the equal-filling approximation Perez-Martin and Robledo (2008) (sHF-EFA) that do not include pairing correlations by construction. Results are indeed very close overall, except in 56,58Ca (0f5/2 shell) where sHFB better reproduces experimental values for $S_{2n}$ and $\Delta_{2n}$. As for the curvature within open-shells, the left panels of Fig. 2 reveals that the curvature of sHF-EFA results also carries the wrong sign but is such that the concavity is even more pronounced than for sHFB, i.e. the weak pairing correlations present within the 0f7/2 shell in sHFB do improve the situation compared to the case where pairing would indeed be strictly zero.
@@
Equation (3a) proves that the sHF-EFA energy is indeed quadratic [^7] As shown in A, the 3N interaction actually induces the presence of a cubic term in the energy. However, present numerical applications demonstrate that it is negligible for all nuclei under consideration such that it can be dropped altogether in the present discussion. in the number of valence nucleons throughout any given open-shell. The coefficient $\alpha_{\breve{v}}$ of the linear term is nothing but the mean-field single-particle energy of the valence shell computed in the CS core $\epsilon^{\text{CS}}_{\breve{v}}$, whose interaction energy contributions are displayed diagrammatically in Fig. 3. The coefficient $\beta_{\breve{v}}$ of the quadratic term, i.e. the curvature of the energy, is given by the average over the valence magnetic substates of the diagonal valence-shell two-body matrix elements [^8] As seen from Eq. (4b), $\overline{\mathbb{v}}_{vv^{\prime}vv^{\prime}}$ includes the effective contribution obtained by averaging the 3N interaction over the CS core. $\overline{\mathbb{v}}_{vv^{\prime}vv^{\prime}}$ displayed diagrammatically in Fig. 4. Such an averaging corresponds to the monopole valence-shell matrix element per valence state. As visible from Eq. (3b), $-2\epsilon^{\text{CS}}_{\breve{v}}$ sets the initial value of $S_{2n}$ [^9] As seen from Tab. 1, the relation $|\alpha_{\breve{v}}|\gg|\beta_{\breve{v}}|$ holds in practice such that the starting value of $S^{\text{sHF-EFA}}_{2n}$ ($a_{v}=2$) in the open shell is essentially dictated by $\epsilon^{\text{CS}}_{\breve{v}}$. whereas $-2\beta_{\breve{v}}$ drives its linear evolution throughout the open-shell. Eventually, $\Delta_{2n}$ extracts $4\beta_{\breve{v}}$.
@@
Figure 5: Neutron canonical single-particle energies $\epsilon_{\breve{k}}$ from sHFB calculations along the Ca isotopic chain. Semi-analytical sHFB-ZP results for $\epsilon_{0\text{f}_{7/2}}$ are also shown between 40Ca and 48Ca.
Extracting $\epsilon^{\text{CS}}_{\breve{v}}$ and $\overline{\mathbb{v}}_{vv^{\prime}vv^{\prime}}$ numerically from the presently employed chiral Hamiltonian (see Tab. 1), the semi-analytical results from Eqs. (3)-(4) (in fact of their sHFB-ZP counterparts; see A) are superimposed on the left panels of Fig. 1 between 42Ca and 48Ca (0f7/2 valence shell) as well as between 62Ca and 70Ca (0g9/2 valence shell). The results perfectly match the numerical sHF-EFA curves, that are themselves very close to sHFB results. Looking at the left panels of Fig. 2, one indeed sees the fully quantitative agreement between sHF-EFA and the semi-analytical results.
@@
    $E^{\text{sHFB}}$ loses energy relatively to experiment,
@@
    $S^{\text{sHFB}}_{2n}$ starts from too low a value,
@@
the coefficient of the slope being given by $\beta_{\tilde{v}}$. As visible in Fig. 5, neutron canonical single-particle energies do evolve linearly within a given open-shell. In particular, the evolution of $\epsilon_{0\text{f}_{7/2}}$ between 40Ca and 48Ca is perfectly reproduced using Eq. (5) (in fact its sHFB-ZP counterpart; see A). Eventually, this linear down-slopping evolution is fully correlated with the concavity of the binding energy.
@@ 3.3 Cr chain
Having characterized sHFB results along the semi-magic Ca isotopic chain, we focus on to doubly open-shell Cr isotopes.
As seen in the upper-right panel of Fig. 1, the global trend of sHFB binding energies is similar, relative to the data, than for Ca isotopes. In the magnifying glass of $S_{2n}$ and $\Delta_{2n}$, experimental data do not however display the characteristic patterns identified along the Ca chain. In particular, $S_{2n}$ decreases more gradually such that the sudden drops (sudden spikes in $\Delta_{2n}$), e.g. at $N=20$ and $28$, have all disappeared. Contrarily, a small bump (spike) is now visible in $S_{2n}$ ($\Delta_{2n}$) for $N=24$, i.e. in 44Cr located in the middle of the 0f7/2 shell. These changes are not at all accounted for by sHFB results that closely follow those obtained previously. Indeed, in addition to displaying the defects identified along the Ca isotopic chain, sHFB results further fail to capture the qualitative modifications seen in the data, i.e. sHFB results keep a strong memory of the underlying spherical shell structure whose fingerprints are no longer visible along the Cr isotopic chain.
@@ 4.1 Ca chain
In spite of the $N=20$ two-neutron shell gap being still overestimated in sBMBPT(2) and sBCCSD, the $S_{2n}$ at the beginning of each open shell is increased to be in much better agreement with experimental data. For example, dynamical correlations bring $S_{2n}$ in 42Ca from $13.3$ MeV in sHFB to $18.0$ and $18.1$ MeV in BMBPT(2) and BCCSD, respectively, as well as to $18.6$ MeV in VS-IMSRG(2), which compares favorably with the experimental value of $19.8$ MeV. Third, the wrong linear increase throughout any given open shell is corrected for, as can be seen for example between 42Ca and 48Ca. This reflects the improvement of the curvature of the energy throughout open shells that can be better appreciated from the left-panels of Fig. 7 that focuses on the 0f7/2 shell. Dynamical correlations turn the energy from beyond concave at the sHFB level to being convex, in a way that is essentially identical with the three employed methods.
Eventually, the agreement with data for $S_{2n}$ and $\Delta_{2n}$ along the Ca chain is qualitatively and quantitatively satisfying for all three methods even though the $N=20$ magicity is still exaggerated in BMBPT(2) and BCCSD and the convexity throughout the 0f7/2 shell is not pronounced enough compared to experimental data for all three methods, which points to yet missing correlations. It will be interesting to investigate in the future whether the lack of convexity in the energy is correlated with the inability of presently employed ab initio methods to correctly reproduce the (infamous) evolution of charge radii between 40Ca and 48Ca Caurier et al. (2001).
@@ 4.2 Analytical investigation
As seen from Eqs. (6)-(8), and in agreement with the results shown in the middle-left panel of Fig. 6 and analyzed in the present section, dynamical correlations modify both the starting value and the slope of $S_{2n}$ in the valence-shell. For example, the negative second-order self-energy correction $\Sigma^{(2)}_{0\text{f}_{7/2}}$ lowers $\epsilon^{\text{CS}(2)}_{0\text{f}_{7/2}}$ in such a way that $S_{2n}$ computed in sBMBPT(2) increases from $13.28$ to $18.02$ MeV in 42Ca to almost match the experimental value ($19.84$ MeV). This effect relates to the coupling of a propagating nucleon to 1-particle/2-hole and 2-particle/1-hole configurations as represented in Fig 8, the latter winning over the former [^11] The lowering of $\epsilon_{0\text{f}_{7/2}}$ is not accompanied by a decrease of $\Delta_{2n}$ in 40Ca in sBMBPT(2) and sBCCSD, contrary to sVS-IMSRG(2), i.e. $\epsilon_{0\text{d}_{3/2}}$ is lowered as much as $\epsilon_{0\text{f}_{7/2}}$. Thus, the needed increase of the effective mass associated with the compression of on-shell single-particle energies is not accounted for by low-order corrections to sHFB.. Consistently, the second-order correction to the average 0f7/2 valence-shell effective interaction is repulsive, with the hole-hole contribution winning over the particle-particle one. In the present calculation, such a correction is larger in absolute value than the mean-field contribution and manages to turn the total energy from being concave to being convex, i.e. it makes $S_{2n}$ decrease linearly between 42Ca and 48Ca as for experimental data [^12] The amount by which $S_{2n}$ is increased at the start of the open-shell and the fact that its slope is actually inverted depend on the Hamiltonian under use; see Refs. Tichai et al. (2018); Somà et al. (2021) for examples where the qualitative defects of the sHFB results are not actually corrected via the inclusion of low-order dynamical correlations.. Still, and as can be seen from the bottom-left panel of Fig. 7, the positive curvature $\beta^{(2)}_{0\text{f}_{7/2}}=25$ keV [^13] This value is essentially constant throughout the valence shell. is not large enough [^14] The same is true for sBCCSD and VS-IMSRG(2) calculations as can be inferred from the bottom-left panel of Fig. 7. compared to experimental data ($\Delta_{2n}/4\approx 220$ keV in 42-46Ca), thus pointing to yet missing many-body correlations as discussed earlier on.